• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name D.Wang et el.

    Log entry time 07:07:23 on November 07, 2009

    Entry number 298038


    keyword=LHRS DAQ problem fixed

    The problem turned out to be a bad connection of L1A signal from the retiming board to the TM module. I unplugged the cable and plugged in back in at the retiming board and got the gates signals back (checking from the scope). And the newly started runs have data now. The pins of the retiming board are kind of loose. If you touch the L1A cable at where it is connected to the board, sometimes you might loose the gate signals.

    A copy of this log entry has been emailed to: rom, rsubedi